OpenStudy (anonymous):

Help! :) Evaluate 6x ÷ 3y - 4, if x = 20 and y = 2.

OpenStudy (konradzuse):

ok so lets first substitute. (6 * 20) / ((3 * 2)-4) =?

OpenStudy (anonymous):

60

OpenStudy (konradzuse):

120/(6-4) =?

OpenStudy (konradzuse):

120/2 =?

OpenStudy (anonymous):

60

OpenStudy (konradzuse):

You guys need to please read the code of conduct, don't just give answers, explain and help out. We askers want to learn, not just solve.

OpenStudy (anonymous):

good

OpenStudy (konradzuse):

No, it's not good, it's horrible.

OpenStudy (konradzuse):

Taylor do you understand?

OpenStudy (anonymous):

Hi @Raja99 and @zibeauty - @KonradZuse is completely correct. Please take a moment to review our Code of Conduct - http://openstudy.com/code-of-conduct when you have a chance. I realize you are trying to help, but we've found that the best way to help students is to try and help them understand the problem, and learn how to solve the problems, not simply provide the answers. Thank you!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Cheers @Raja99 :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Thanks @zibeauty - much appreciated ;)